# SpikeDet: Better Firing Patterns for Accurate and Energy-Efficient Object Detection with Spiking Neuron Networks

### v3

@@ V-C Comparisons with State-of-The-Art Models
V-D Evaluation on Object Detection Sub-Tasks
We further apply our proposed SpikeDet to two object detection sub-tasks, event camera object detection on the GEN1 dataset [27] and underwater object detection on the URPC 2019 dataset [28], and compare it with ANN and SNN baseline methods in these domains. Implementation details and dataset introductions are presented in Appendix B.
@@ V-D Evaluation on Object Detection Sub-Tasks
V-D1 Event Camera Object Detection
As shown in Table VI, on the GEN1 dataset, our method not only significantly outperforms SNN-based object detectors but also surpasses all ANN-based methods while maintaining the energy efficiency advantage of SNNs. This marks the first time an SNN-based method surpasses ANNs, demonstrating that firing patterns are more crucial for SNNs in processing event data, enabling better alignment between them. Notably, our model achieves only a 1.1% AP drop when T×$\times$×D is 5$\times$1 compared to 4$\times$2, while significantly reducing firing rate and power consumption. This demonstrates that improved firing patterns can enable SNN models to achieve excellent performance using only LIF neurons when processing event data. Furthermore, we replace I-LIF neurons with ReLU to implement an ANN version of SpikeDet on GEN1. This results in substantial performance degradation and significantly increased power consumption, highlighting the advantages of SNNs in processing event data.
@@ V-D1 Event Camera Object Detection
Model AP AP50 Param (M) T×$\times$×D Firing Rate(%) Power (mJ)
EfficientDet-D3 [59] - 74.8 12.0 1 - 7.34
YOLOv5-S [67] 42.0 75.3 7.2 1 - 9.5
YOLOv8-N [38] 43.6 77.7 3.2 1 - 5.0
ULO [28] - 65.1 3.8 1 - 3.0
YOLOv9-S-UI [68] 48.7 78.1 4.1 1 - 10.7
Spiking-YOLO [13] 22.4 53.7 8.7 3500 - -
SU-YOLO [47] 42.9 78.8 7.0 4 - 3.0
SpikeYOLO-N [20] 46.9 83.1 14.2 1$\times$4 12.5 5.8
SpikeYOLO-S [20] 47.9 83.8 23.1 1$\times$4 12.2 8.6
SpikeDet-S 49.2 84.3 22.0 1$\times$4 11.2 5.5
@@
V-D2 Underwater Object Detection
@@
As shown in Table VII, on the URPC 2019 dataset, SpikeDet achieves significant improvements compared to other methods in both accuracy and power consumption, reaching state-of-the-art performance among all methods. Specifically, compared to YOLOv9-S-UI, it achieves 1.4% AP improvement while reducing power consumption by 5.2 mJ.
VI Discussion on Neuromorphic Chip Deployment
In the above sections, we have conducted theoretical evaluations to compare power consumption between ANNs and SNNs, following common practices in SNN research [17, 15, 47, 19, 21]. It is worth noting that the theoretical analysis excludes hardware architecture or data caching. Efficiency gains may be further amplified when SNNs are deployed on neuromorphic chips. These neuromorphic platforms, such as the Speck chip [12], operate in asynchronous mode without a global clock, leveraging spike-driven characteristics. Since I-LIF neurons are naturally suited for asynchronous operations [21], our approach can fully exploit this spike-driven paradigm. This design leads to minimal static power consumption during periods without spike inputs. Consequently, the overall energy consumption of neuromorphic computing primarily scales with the number of spike-induced ACs, further enhancing the advantages of SNNs over ANNs when static power consumption is considered.
For deployment to real hardware, we take the chip Speck as an example. First, we need to use Sinabs, a PyTorch-based deep learning library, to build and train SNN models. Through the Sinabs-Speck plugin, we can complete model quantization and configuration generation. Subsequently, we can use the Samna tool for deployment and execution.
VII Conclusion
@@
In this paper, to overcome the local firing saturation problem and achieve better accuracy and energy efficiency, we propose SpikeDet, a novel SNN-based object detector. Specifically, the key contribution of our approach is MDSNet, which employs MDS to enhance membrane-based shortcuts with output distribution adjustment capabilities, thereby stabilizing the synaptic input distribution of neurons and achieving better neuron firing patterns in backbone features. Additionally, for the feature fusion module, we propose SMFM, which performs multiple refinement processes through a multi-directional fusion strategy, enhancing the multi-scale detection capability while preserving firing patterns. Experimental results demonstrate that our model outperforms other SNN-based models in accuracy on the COCO 2017 dataset, and even surpasses ANN-based methods on downstream tasks such as GEN1 and URPC 2019, proving its strong generalization capability. Meanwhile, it enhances SNN advantages in power consumption, achieving a superior performance-efficiency trade-off. In the future, we plan to investigate whether similar local firing saturation issues exist in other SNN-based vision tasks and propose corresponding solutions.
@@ A-A Proof of Proposition 2
Proposition 2. For I-LIF neurons, if the input follows $\mathbf{x}^{t,n-1}\sim\mathcal{N}(0,{\sigma_{in}}^{2})$, then the membrane potential follows $\mathbf{u}^{t,n}\sim\mathcal{N}(0,{\sigma_{out}}^{2})$, where ${\sigma_{out}}^{2}$ is proportional to ${\sigma_{in}}^{2}$.
@@
$=\tau[\tau(\mathbf{u}^{t-2,n}-\mathbf{o}^{t-2,n})-\mathbf{o}^{t-1,n}]+\mathbf{x}^{t,n-1}$
$={\tau}^{t-1}\mathbf{x}^{1,n-1}-\sum_{p=1}^{t-1}{\tau}^{p}\mathbf{o}^{t-p,n}+\sum_{p=1}^{t}{\tau}^{p}\mathbf{x}^{t-p,n-1}.$
Since $\tau$ is a tiny constant, which is only 0.25 in our implementation, the above equation can be simplified to:
$\mathbf{u}^{t,n}=\mathbf{x}^{t,n-1}+{\tau}\mathbf{x}^{t-1,n-1}-\tau\mathbf{o}^{t-1,n}.$ (16)
Moreover, since I-LIF can fire integer-valued spikes, each spike emission reduces the membrane potential to a level close to 0. Therefore, we can assume that the last spike is fired at time step $t^{\prime}$, and thus ignore the influence of $\mathbf{o}$ on the membrane potential, yielding:
@@
Because we employ tdBN for normalization, x is independently and identically distributed for each time step $t$. Thus, we can conclude that $\mathbf{u}^{t,n}\sim N(0,{\sigma_{out}}^{2})$, where ${\sigma_{out}}^{2}$ is proportional to ${\sigma_{in}}^{2}$.
TABLE VIII: The Structure of MDSNet for COCO 2017. The Output Size column specifies the spatial resolution and channel count of feature maps produced at each stage. The numbers in parentheses indicate the number of MS-Blocks contained in the residual path of the MDS-Block.
@@
Encoding 320$\times$320, 64 Encoding Block
Conv1 160$\times$160, 64 MS2Block2(0) MS2Block2(1) MS2Block2(2) MS2Block2(2)
Conv2 80$\times$80, 128 MS2Block2(0) MS2Block2(1) MS2Block2(1) MS2Block1(1) MS2Block2(3) MS2Block1(3)
Conv2 40$\times$40, 256 MS2Block2(0) MS2Block2(1) MS2Block2(2) MS2Block1(2) MS2Block2(15) MS2Block1(15)
Conv4 20$\times$20, 512 MS2Block2(0) MS2Block2(1) MS2Block2(2) MS2Block2(3) MS2Block1(3)
Conv5 20$\times$20, 512 SpikeSPPF
A-B Proof of Proposition 3
@@
Theorem 2. (Theorem 4.2. in [50]) Given $\mathbf{J}:=\sum_{j}\mathbf{J_{j}}$, where ${\mathbf{J_{j}}}$ is a series of independent random matrices. If at most one matrix in ${\mathbf{J_{j}}}$ is not a central matrix, we have
@@
For the residual path, since all its components satisfy Definition 2 for MDS-Block1 and MDS-Block2 [50, 15, 33], we can derive:
@@ B-1 Datasets Introduction
The GEN1 dataset [27] represents the initial large-scale collection for object detection using event cameras. It comprises car footage spanning over 39 hours, captured by the GEN1 device with a spatial resolution of 304$\times$240. The dataset includes bounding box annotations for vehicles and pedestrians, provided at rates of 1 to 4Hz. These labels accumulate to a total exceeding 255,000 across the recordings.
The URPC 2019 dataset [28] is designed for the 2019 Underwater Robot Picking Contest and comprises 4,707 annotated underwater images across four marine object categories: scallops, starfish, echinus, and holothurian. The dataset splits include 3,767 training images, 695 validation images, and 245 test images.

### v4

@@ V-D Comparisons with State-of-The-Art Models
V-E Evaluation on Object Detection Sub-Tasks
We further apply our proposed SpikeDet to four object detection sub-tasks, event camera on the GEN1 dataset [tournemire_large_2020], underwater on the URPC 2019 dataset [wang_ulo_2022], low-light on the ExDARK dataset [loh2019getting], and dense scene on the CrowdHuman dataset [shao2018crowdhuman]. These challenging scenarios allow us to evaluate the generalization capability and robustness of our method across diverse conditions. Implementation details and dataset introductions are presented in Appendix B.
TABLE VI: Performance Comparison with State-of-the-Art Models on GEN1 Dataset. The -ReLU indicates that the activation function of SpikeDet-S is changed to ReLU, converting it to an ANN version.
[…]
RED [perot_learning_2020] 40.0 - 24.1 - - - >>24.1
SpikeDet-S-ReLU 46.2 66.5 22.0 1 - - 45.2
ASTMNet [perot_learning_2020] 46.7 - >>100 3 - - -
RVT [gehrig_recurrent_2023] 47.2 - 18.5 10 - - 116.7
S5-Vit-B [zubic_state_2024] 47.4 - 18.2 10 - - -
VC-DenseNet [cordone_object_2022] 18.9 - 24.3 5 37.2 - 5.8
Spiking-YOLO [kim_spikingyolo_2020] - 45.3 7.9 5 - - 0.9
EMS-YOLO [su_deep_2023] 31.0 59.0 14.4 5 17.8 - 3.4
SFOD [fan_sfod_2024] 32.1 - 11.9 5 24.4 7.5 7.3
EAS-SNN [wang_eassnn_2025] 37.5 69.9 25.3 5 20.3 - 28.1
SpikingViT-B [yu_spikingvit_2024] 39.4 61.6 21.5 >>64 18.0 - -
[…]
V-E1 Event Camera Object Detection
[…]
@@ V-E1 Event Camera Object Detection
EfficientDet-D3 [tan_efficientdet_2020] - 74.8 12.0 1 - - 7.3
YOLOv8-N [yolov8_ultralytics] 43.6 77.7 3.2 1 - - 5.0
ULO [wang_ulo_2022] - 65.1 3.8 1 - - 3.0
YOLOv9-S-UI [pan_optimization_2024] 48.7 78.1 4.1 1 - - 10.7
Spiking-YOLO [kim_spikingyolo_2020] 22.4 53.7 8.7 3500 - - -
SU-YOLO [li_suyolo_2025] 42.9 78.8 7.0 4 - - 3.0
[…]
V-E2 Underwater Object Detection
@@
As shown in Table VII, on the URPC 2019 dataset, SpikeDet achieves significant improvements compared to other methods in both accuracy and power consumption. Specifically, compared to YOLOv9-S-UI, it achieves 1.4% AP improvement while reducing power consumption by 5.7 mJ. The superior performance of SpikeDet can be attributed to its more substantial LFSI reduction on URPC 2019 compared to other SNN-based methods. This is particularly beneficial for underwater scenarios where low-contrast and blurred objects demand enhanced feature discrimination for accurate detection.
TABLE VIII: Performance Comparison with State-of-the-Art Models on ExDARK Dataset. $\dagger$ indicates results from multiscale testing.
[…]
DETR-R50 [carion_endtoend_2020] 25.0 55.0 41.3 1 - - 219.0
YOLOv8-S [yolov8_ultralytics] 40.7 66.2 11.1 1 - - 65.6
[…]
V-E3 Low-Light Object Detection
As shown in Table VIII, our method demonstrates robust performance under low-light conditions. By effectively mitigating local firing saturation, SpikeDet achieves a 2.6% AP improvement and 18.9 mJ power reduction compared to SpikeYOLO. Meanwhile, it also exhibits a better performance-power trade-off compared to ANN-based methods.
TABLE IX: Performance Comparison with State-of-the-Art Models on CrowdHuman Dataset. $\dagger$ indicates results from multiscale testing.
[…]
DETR-R50 [carion_endtoend_2020] 27.8 55.5 41.3 1 - - 219.0
YOLOv8-S [yolov8_ultralytics] 51.7 79.3 11.1 1 - - 65.3
[…]
V-E4 Dense Scene Object Detection
As shown in Table IX, SpikeDet achieves strong performance on the CrowdHuman dataset. It outperforms SpikeYOLO by 4.3% AP while consuming 19.2 mJ less power. In such crowded scenarios with heavy occlusions, the lower LFSI enables better discrimination of overlapping objects, contributing to improved detection accuracy.
VI Conclusion
@@
In this paper, we identify the local firing saturation problem in existing SNN-based object detection methods. This issue limits the achievement of better accuracy and energy efficiency. To address this, we propose SpikeDet, a novel SNN-based object detector. The key contribution is MDSNet, which effectively stabilizes synaptic input distribution and mitigates local firing saturation through the proposed MDS and architectural improvements. For feature fusion, we propose SMFM, which employs a multi-directional fusion strategy to enhance multi-scale detection while preserving neuron firing patterns. Additionally, we introduce the Local Firing Saturation Index (LFSI) to quantify the effectiveness of our approach in addressing local firing saturation. Experimental results demonstrate that our model outperforms other SNN-based models in accuracy on the COCO 2017 dataset. Moreover, it achieves superior performance on diverse downstream tasks, including event-based GEN1, underwater URPC 2019, low-light ExDARK, and dense scene CrowdHuman, demonstrating strong generalization capability. Meanwhile, it maintains significantly lower power consumption, achieving an optimal balance between detection accuracy and energy efficiency.
@@ A-A Proof of Proposition 1
Proposition 1. For $d$ stacked LCB layers ($d\geq 1$) with arbitrary kernel sizes, under the convolution zero-mean weight and zero bias initialization, the output at the $(n+d)$-th layer $\mathbf{y}^{t,n+d}$ is uncorrelated with the input at the $n$-th layer $\mathbf{x}^{t,n}$.
We first analyze the case of a single LCB, conducting our analysis according to the processing sequence of its internal components. The first component is the I-LIF neuron. It receives input $\mathbf{x}^{t,n}$ and produces output spikes $\mathbf{o}^{t,n}=\mathit{Clip}\left(\mathit{round}(\mathbf{u}^{t,n}),0,D\right)$.
Subsequently, we combine the convolution with tdBN following [zheng_going_2021], where the merged weight and bias are:
$W^{\prime}_{m,n}=\lambda_{n}\frac{\alpha V_{th}W_{m,n}}{\sqrt{\sigma^{2}_{inf,n}+\epsilon}},$ (22)
$B^{\prime}_{m,n}=\lambda_{n}\frac{\alpha V_{th}(B_{m,n}-\mu_{inf,n})}{\sqrt{\sigma^{2}_{inf,n}+\epsilon}}+\beta_{n}.$ (23)
Here $W$ represents the weight and $B$ represents the bias, with subscripts $m$, $n$ respectively indicating the mapping from the $m$-th feature map of the previous layer to the $n$-th feature map of the next layer, where the superscript $\prime$ indicates the transformed version. For the parameters in tdBN, $\lambda_{n}$ and $\beta_{n}$ are learnable parameters, $\alpha$ is a hyperparameter, and $\sigma^{2}_{inf,n}$ and $\mu_{inf,n}$ are the expectations of $E[\mathbf{x^{t,n}}]$ and $Var[\mathbf{x^{t,n}}]$ over the entire dataset.
Taking the expectation of the merged weight:
$E[W^{\prime}_{m,n}]=\lambda_{n}\frac{\alpha V_{th}}{\sqrt{\sigma^{2}_{inf,n}+\epsilon}}E[W_{m,n}].$ (24)
Since $E[W_{m,n}]=0$ (zero-mean weight initialization) and $\lambda_{n},\alpha,V_{th},\sigma^{2}_{inf,n},\epsilon$ are deterministic constants:
$E[W^{\prime}_{m,n}]=0.$ (25)
Thus, tdBN merging preserves the zero-mean property of weights. Furthermore, given that $B_{m,n}$ is initialized to zero, and according to the default parameters in [zheng_going_2021], $\mu_{inf,n}$ and $\beta_{n}$ are also 0, so from Eq. 23, we get $B^{\prime}_{m,n}=\mathbf{0}$.
As in [he2015delving, glorot2010understanding, taki_deep_2017], we assume that the elements in $\mathbf{x}^{t,n}$ and $W_{m,n}$ are independent and identically distributed (i.i.d.). For rigorous covariance analysis, we consider individual tensor elements. Let $y_{i}^{t,n}$, $x_{i}^{t,n}$, $w_{i}^{t,n}$, $b_{i}^{t,n}$, and $o_{i}^{t,n}$ denote arbitrary elements of $\mathbf{y}^{t,n}$, $\mathbf{x}^{t,n}$, $W^{\prime}_{m,n}$, $B^{\prime}_{m,n}$, and $\mathbf{o}^{t,n}$, respectively. The $\mathbf{y}^{t,n}$ represents the output of a single LCB, and its element $y_{i}$ can be expressed as
$y_{i}^{t,n}=\sum_{j\in\text{RF}(i)}w_{j}^{n}\cdot o_{j}^{n}+b_{i}^{t,n},$ (26)
where $\text{RF}(i)$ denotes the receptive field corresponding to output position $i$. For simplicity,
$y_{i}^{t,n}=\sum_{j}w_{j}^{t,n}\cdot o_{j}^{t,n}+b_{i}^{t,n}.$ (27)
From the above derivation, we can obtain $b_{i}^{t,n}=0$, therefore,
$y_{i}^{t,n}=\sum_{j}w_{j}^{t,n}\cdot o_{j}^{t,n}.$ (28)
Then, we derive the covariance between the output $y_{i}^{t,n}$ and the input $x_{i}^{t,n}$,
$\text{Cov}(y_{i}^{t,n},x_{i}^{t,n})=E[y_{i}^{t,n}\cdot x_{i}^{t,n}]-E[y_{i}^{t,n}]E[x_{i}^{t,n}].$ (29)
Since $E[x_{i}^{t,n}]=0$,
$\text{Cov}(y_{i}^{t,n},x_{i}^{t,n})=E[y_{i}^{t,n}\cdot x_{i}^{t,n}].$ (30)
Expanding the above equation, we get
$E[y_{i}^{t,n}\cdot x_{i}^{t,n}]=E\left[\left(\sum_{j}w_{j}^{t,n}\cdot o_{j}^{t,n}\right)\cdot x_{i}^{t,n}\right].$ (31)
Exchanging summation and expectation,
$E[y_{i}^{t,n}\cdot x_{i}^{t,n}]=\sum_{j}E[w_{j}^{t,n}\cdot o_{j}^{t,n}\cdot x_{i}^{t,n}].$ (32)
Since weights $w_{j}^{t,n}$ are independent of inputs at initialization,
$E[w_{j}^{t,n}\cdot o_{j}^{t,n}\cdot x_{i}^{t,n}]=E[w_{j}^{t,n}]\cdot E[o_{j}^{t,n}\cdot x_{i}^{t,n}].$ (33)
Therefore:
$E[w_{j}^{t,n}\cdot o_{j}^{t,n}\cdot x_{i}^{t,n}]=0\cdot E[o_{j}^{t,n}\cdot x_{i}^{t,n}]=0.$ (34)
This holds for all $j$. From Eq. 32:
$E[y_{i}^{t,n}\cdot x_{i}^{t,n}]=\sum_{j}E[w_{j}^{t,n}\cdot o_{j}^{t,n}\cdot x_{i}^{t,n}]=\sum_{j}0=0.$ (35)
$\text{Cov}(y_{i}^{t,n},x_{i}^{t,n})=0.$ (36)
Therefore, for a single LCB, its output $\mathbf{y}$ is uncorrelated with the input $\mathbf{x}$.
Finally, we consider $d$ stacked LCBs ($d\geq 1$) and analyze them using mathematical induction. Assume that for (d1)(d-1) stacked LCBs, we have $\text{Cov}(y^{t,n+d-1},x^{t,n})=0$. We prove that this also holds for $d$ stacked LCBs, i.e., $\text{Cov}(y^{t,n+d},x^{t,n})=0$.
The output of layer $(n+d)$ is:
$y^{t,n+d}_{i}=\sum_{j}w^{t,n+d}_{j}\cdot o^{t,n+d}_{j}.$ (37)
Since $E[\mathbf{x}^{t,n}]=0$,
$\text{Cov}(y^{t,n+d}_{i},x^{t,n}_{i})$ $=E[y^{t,n+d}_{i}\cdot x^{t,n}_{i}]$ (38)
$=\sum_{j}E[w^{t,n+d}_{j}\cdot o^{t,n+d}_{j}\cdot x^{t,n}_{i}].$
At initialization, weights $w^{t,n+d}_{j}$ are independent of the input $\mathbf{x}^{t,n}$ at layer $n$. Although $o^{t,n+d}_{j}$ depends on $\mathbf{x}^{t,n}$ through intermediate layers, the weight independence holds:
$E[w^{t,n+d}_{j}\cdot o^{t,n+d}_{j}\cdot x^{t,n}_{i}]=E[w^{t,n+d}_{j}]\cdot E[o^{t,n+d}_{j}\cdot x^{t,n}_{i}].$ (39)
Since $E[w^{t,n+d}_{j}]=0$ by zero-mean initialization,
$E[w^{t,n+d}_{j}\cdot o^{t,n+d}_{j}\cdot x^{t,n}_{i}]=0.$ (40)
Thus $\text{Cov}(y^{t,n+d}_{i},x^{t,n}_{i})=0$. By mathematical induction, $\text{Cov}(\mathbf{y}^{t,n+d},\mathbf{x}^{t,n})=0$ for any starting layer $n$ and $d$ stacked LCB layers. This completes the proof of Proposition 2.
A-B Proof of Proposition 2
Proposition 2. For I-LIF neurons, the variance of the membrane potential is proportional to the variance of the input, i.e., $\text{Var}[\mathbf{u}^{t,n}]\propto\text{Var}[\mathbf{x}^{t,n}]$.
@@ A-B Proof of Proposition 2
$=\tau[\tau(\mathbf{u}^{t-2,n}-\mathbf{o}^{t-2,n})+\mathbf{x}^{t-1,n-1}-\mathbf{o}^{t-1,n}]+\mathbf{x}^{t,n-1}$
$={\tau}^{t}\mathbf{u}^{0,n}-\sum_{p=1}^{t-1}{\tau}^{t-p}\mathbf{o}^{p,n}+\sum_{p=1}^{t}{\tau}^{t-p}\mathbf{x}^{p,n-1}.$
Since I-LIF can fire integer-valued spikes, each spike emission reduces the membrane potential to a level close to 0. Therefore, we can assume that the last spike is fired at time step $t^{\prime}$, and thus ignore the influence of $\mathbf{o}$ on the membrane potential, yielding:
$\mathbf{u}^{t,n}={\tau}^{t}\mathbf{u}^{0,n}+\sum_{p=t^{\prime}}^{t}{\tau}^{t-p}\mathbf{x}^{p,n-1}.$ (42)
Moreover, since $\tau$ is a tiny constant, which is only 0.25 in our implementation, the above equation can be simplified to:
@@
Assume that $\mathbf{x}$ is independently and identically distributed for each time step $t$. Thus, we can conclude that the membrane potential variance is proportional to the input variance, i.e., $\text{Var}[\mathbf{u}^{t,n}]\propto\text{Var}[\mathbf{x}^{t,n-1}]$.
TABLE X: The Structure of MDSNet for COCO 2017. The Output Size column specifies the spatial resolution and channel count of feature maps produced at each stage. The numbers in parentheses indicate the number of MS-Blocks contained in the residual path of the MDS-Block.
@@
Encoding $\times320$, 64 Encoding Block
Conv1 $\times160$, 64 MS2Block2(0) MS2Block2(1) MS2Block2(2) MS2Block2(2)
Conv2 $\times80$, 128 MS2Block2(0) MS2Block2(1) MS2Block2(1) MS2Block1(1) MS2Block2(3) MS2Block1(3)
Conv2 $\times40$, 256 MS2Block2(0) MS2Block2(1) MS2Block2(2) MS2Block1(2) MS2Block2(15) MS2Block1(15)
Conv4 $\times20$, 512 MS2Block2(0) MS2Block2(1) MS2Block2(2) MS2Block2(3) MS2Block1(3)
Conv5 $\times20$, 512 SpikeSPPF
A-C Proof of Proposition 3
@@
Theorem 2. (Theorem 4.2. in [chen_comprehensive_2020]) Given $\mathbf{J}:=\sum_{j}\mathbf{J_{j}}$, where $\mathbf{J_{j}}$} is a series of independent random matrices. If at most one matrix in $\mathbf{J_{j}}$} is not a central matrix, we have
@@
For the residual path, since all its components satisfy Definition 2 for MDS-Block1 and MDS-Block2 [chen_comprehensive_2020, su_deep_2023, zheng_going_2021], we can derive:
@@ B-1 Datasets Introduction
[…]
The URPC 2019 dataset [wang_ulo_2022] is designed for the 2019 Underwater Robot Picking Contest and comprises 4,707 annotated underwater images across four marine object categories: scallops, starfish, echinus, and holothurian. The dataset splits include 3,767 training images, 695 validation images, and 245 test images.
The ExDARK dataset [loh2019getting] is designed for low-light object detection, featuring 7,363 images captured under various illumination conditions ranging from extremely dark environments to twilight scenarios. The dataset encompasses 12 object categories: Bicycle, Boat, Bottle, Bus, Car, Cat, Chair, Cup, Dog, Motorbike, People, and Table, with bounding box annotations provided for all instances. In this study, we split the dataset into training, validation, and test sets with a ratio of 6:2:2.
The CrowdHuman dataset [shao2018crowdhuman] focuses on human detection in densely populated scenarios. It comprises approximately 470K annotated human instances across its training and validation sets, with an average density of 22.6 individuals per image and substantial occlusion variations. Annotations include bounding boxes for the head, visible body region, and complete body for each person. The dataset provides 15,000 training images, 4,370 validation images, and 5,000 test images, featuring comprehensive annotations across diverse scenarios.
Figure 12: Detection result on the URPC 2019 dataset.
Figure 13: Detection result on the ExDARK dataset.
Figure 14: Detection result on the CrowdHuman dataset.